OpenStudy (rachel1234):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5 I had no idea what I'm doing so I'd prefer just an easy answer. I need to get this stuff done with

Nnesha (nnesha):

\[\huge\rm 7\sin^2x-14sinx+2=-5\] set it equal to 0 (add 5 both sides) \[\large\rm 7\sin^2x-14sinx+2+5=0\]\[\large\rm 7\sin^2x-14sinx+7=0\] this is just a quadratic equation let say `sinx=y`\[\large\rm 7y^2-14y+7=0\] find the solution of this quadratic equation

OpenStudy (rachel1234):

I don't even know how to do that

Nnesha (nnesha):

you don't know how to solve quadratic equation ?

Nnesha (nnesha):

take out the common factor which is 7\[7(y^2-2y+1)=0\] \[\large\rm y^2-2y+1=0\] you can use quadratic formula \[\huge\rm \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] where a=leading coefficient b=middle term( coefficient of `x` term) C= constant \[\large\rm Ax^2+Bx+C=0\]

Nnesha (nnesha):

7(y^2-2y+1)=0 divide both sides by 7. that's how i got y^2-2y+1=0